OpenStudy (anonymous):

Two is thirty-one, divided by the sum of a number and four. Find the number. A. 11.5 B. –15.5 C. 13.5 D. none of the above

OpenStudy (anonymous):

2 = 31/n+4 2n+8 = 31 2n = 23 n = 11.5

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

tim spent$24 at the grocery store. This was six dollars less than two times what he spent at the game store. how much did he spend at the game store

OpenStudy (anonymous):

2g -6 = 24 2g = 30 g = 15

OpenStudy (anonymous):

when 5 is added 2 times a number, the result is 45. find the number

OpenStudy (anonymous):

2n + 5 = 45 2n = 40 n = 20

OpenStudy (anonymous):

Ken has a pole that is 26 inches long. he wishes to cut it into two pieces so that one piece will be 8 inches longer than the other. how long must the shorter piece be?